OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 16.

OpenStudy (anonymous):

\[x^2/64+y^2/81=1\] \[x^2/9+y^2/8=1\] \[x^2/81+y^2/64=1\] \[x^2/8+y^2/9=1\]